nikkixmarie:

A quadratic function was translated 3 up. Give the equation and graph

nikkixmarie:

i have the equation (x-1)^2+k is it right?

nikkixmarie:

@loveboy

loveboy:

Why (x-1)^2 ?

nikkixmarie:

the equation for quadric is (x-h)^2+k but i dont know what to put for h so i put 1

loveboy:

h could be 0 as well and k here would be 3

loveboy:

x-1 denotes you're shifting right to 1 units.

nikkixmarie:

so how would you write it if h is 0

loveboy:

\[\large \bf y=x^2+3\]

nikkixmarie:

ohhh ok if i write the equation for 2 more could you check it?

loveboy:

Sure

nikkixmarie:

8. A square root function was translated 2 left,5 up and was reflected over the x-axis.

nikkixmarie:

i got \[\sqrt{x-2}+5\]

loveboy:

Wrong :(

nikkixmarie:

tbh i dont even know the proper equation for this

loveboy:

First, they asked 2 left, it'd be sqrt(x+2)

loveboy:

And reflection over x-axis, so \[\large \bf y=-\sqrt{(x+2)}+5\]

nikkixmarie:

ohh ok and then when its not reflected you would take out the negative?

loveboy:

Remember this :

loveboy:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @nikkixmarie ohh ok and then when its not reflected you would take out the negative? \(\color{#0cbb34}{\text{End of Quote}}\) Right!

nikkixmarie:

ok thank you literally 2 more to check and ill leave you be! ive been stuck on this topic

loveboy:

Sure

nikkixmarie:

9. A linear function was translated 3 down and reflected. Give the equation and graph equation: -3x+1